THE ELECTRONIC SPECTRA OF THE JET-COOLED BORON DIFLUORIDE (BF_{2}) AND BORON DICHLORIDE (BCl_{2}) FREE RADICALS

\underline{JIE YANG}, JIE WEI, FUMIE SUNAHORI, MOHAMMED GHARAIBEH, BLAINE ELLIS, AL-FRED CHEN, and DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055, USA; SHENG-GUI HE, Beijing National Laboratory for Molecular Sciences, State Key Laboratory for Structural Chemistry of Unstable and Stable Species, Institute of Chemistry, The Chinese Academy of Sciences, Zhongguancun, Beijing 100080, PRC.

Laser-induced fluorescence (LIF) and single vibronic level (SVL) emission spectra of the Ã^{2}Π_{u} − X̃^{2}A_{1} transitions of the jet-cooled boron difluoride (BF_{2}) and boron dichloride (BCl_{2}) radicals have been observed. In each case, the radical was produced in a pulsed discharge jet using a precursor mixture of the corresponding boron halide (BF_{3} or BCl_{3}) in high pressure argon. Due to a large change in geometry from a bent ground state to a linear excited state, the LIF spectra are very extensive and the emission spectra show long bending progressions. Strong LIF bands of ^{11}BF_{2} and ^{10}BF_{2} (from isotopically enriched ^{10}BF_{3}) were observed in the 29200 − 35100 cm^{−1} region and isotope shifts were identified. The rotational constants obtained from high resolution LIF spectra of the two isotopologues were used to determine the effective zero-point structure of the ground state. For BCl_{2}, weaker and much more complex LIF spectra, which have as yet defied vibrational analysis, were recorded in the 15300−22700 cm^{−1} region. SVL emission spectra of five isotopologues have been obtained, affording a detailed vibrational analysis of the ground state of BCl_{2}.